OpenStudy (anonymous):

Gravel is being dumped from a conveyor belt at a rate of 20 cubic feet per minute. It forms a pile in the shape of a right circular cone whose base diameter and height are always equal. How fast is the height of the pile increasing when the pile is 14 feet high? Recall that the volume of a right circular cone with height h and radius of the base r is given by V=1/3πr^2h Rate of change of height = feet per minute

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

iv finished most of it, but now im not sure where to go with it

OpenStudy (anonymous):

this is what i got so far (1/3)pi(2r((1/2)dh/dt)h+r^2dh/dt)

zepdrix (zepdrix):

No you don't want to take the derivative while you have 2 variable in there. That's no good.

OpenStudy (anonymous):

well thats what we did in class, so we can start over if thats easier for you

zepdrix (zepdrix):

Either we're missing some information, or I'm just misunderstanding something perhaps. If it's a `right circular` cone... does that ... oh oh oh they tell us that the diameter and height are equal, ok that's what I was missing.

OpenStudy (anonymous):

brb

zepdrix (zepdrix):

\[\large V=\frac{1}{3}\pi r^2 h\] \[\large r=\frac{1}{2}d=\frac{1}{2}h\]We can determine this since the height and diameter are equal. Substituting in,\[\large V=\frac{1}{3}\pi\left(\frac{1}{2}h\right)^2h\]

zepdrix (zepdrix):

Unlike the last problem, this time, let's make sure we simplify it down before taking a derivative.

zepdrix (zepdrix):

We get something like this, \[\large V=\frac{1}{12}\pi h^3\] Take a look at those steps, make sure they make sense to you.

OpenStudy (anonymous):

yup that makes sense

zepdrix (zepdrix):

The givens are:\[\large V'=20\]\[\large h=14\] Take a derivative, then solve for \(\large h'\)

zepdrix (zepdrix):

It will be just a tad different than the previous problem since you'll need to isolate the h' after differentiating.

OpenStudy (anonymous):

V=1/12πh^3 so dV/dt=1/4pih^2

zepdrix (zepdrix):

You are missing that important term again. :O

OpenStudy (anonymous):

sorry which was?

zepdrix (zepdrix):

dV/dt=1/4pih^2 dh/dt

OpenStudy (anonymous):

oh ya lol my bad

OpenStudy (anonymous):

so dh/dt=20/(1/4pi(14)^2)

zepdrix (zepdrix):

ya looks good.

OpenStudy (anonymous):

thank you so much

zepdrix (zepdrix):

no probs c:

OpenStudy (anonymous):

now hopefully i can pass my test tomorrow for once lol

zepdrix (zepdrix):

heh

OpenStudy (anonymous):

have a great night